What fraction of 4 2/3 is 4/9?
Found 2 solutions by checkley77, josmiceli:
Answer by checkley77(12844)

(4/9)/(4 2/3)
(4/9)/(14/3)
(4/9)*(3/14)
12/126
2/21 is the fractional answer.
proof
14/3*2/21
28/63=4/9

Answer by josmiceli(19441)

First convert 4 2/3 to 1/3 's
There are [3] 1/3's in 1
There are [3*4] 1/3's in 4
So, 4 2/3 = 12/3 + 2/3
12/3 + 2/3 = 14/3
Now the question becomes, What fraction of 14/3 is 4/9?
Now convert 14/3 to 1/9's. Multiply top and bottom by 3
14/3 * 3/3 = 42/9
Now the question becomes, What fraction of 42/9 is 4/9?
(4/9) / (42/9) = 4/42
4/42 = 2/21 answer
You can check by converting to decimals
4 2/3 = 4.6666...
4/9 = .4444...
.4444/4.6666 = .09523
2/21 = .09523
OK